Marion’s grandfather’s will established a trust that will pay her $3200 every three months for 11 years. The first payment will be made six years from now, when she turns 19. If money is worth 8.2% compounded quarterly, what is today’s economic value of the bequest? (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

We need to find the value of annuity payments and calculate its value at present;

PV of annuity due payments= A*(1+r)*[1-(1+r)^-n]/r

where A=annuity payment ; r= interest rate per 3 months =8.2/4 =2.05% ; n= no of quarters= 11*4=44

PV of annuity due=3200*(1+2.05%)*[1-(1+2.05%)^-44]/.0205 =94069.16

This value is of 6 years from now which has to be discounted to present value to find the present worth.

Present value = 94069.16/(1+2.05%)^24 =57800.96